# RAVEN: Rethinking Adversarial Video Generation with Efficient Tri-plane Networks

### v1

@@ 3.3 CNN backbone of Tri-plane and generation
The tri-plane’s composition of three 2D feature arrays allows for the use of a conventional 2D CNN generator. We selected the StyleGAN-T architecture for its exceptional capacity for feature representation. Adapting StyleGAN-T’s output, our model generates a $64\times 64\times 68$ feature grid, partitioned into a 32-channel global feature grid ($F_{G}$) and three $12$-channel motion feature grids. From these, we derive local and global flow fields and a masking channel, which is utilized to warp and blend the appearance features, forming an appearance feature volume. To farther enhance training efficiency, we sample a frame at a random time $t$ at a $64\times 64$ resolution and generate the entire video at $32\times 32\times 32$. We achieve the final $256\times 256$ resolution through image-domain super-resolution, applied only to the selected frame to conserve computational resources. This process mirrors the approach taken in eg3D, utilizing two StyleGAN2 synthesis blocks to upscale the image by a factor of four.
@@ 3.4 Double Discrimination
Similar to a standard GAN setting, we use a discriminator to criticize the generated data. We deploy two different discriminators, one for the frame of time step $t$, and the other for the generated video. We extend the built-in StyleGAN-T discriminator to accommodate $6$ channels instead of $3$. Such that it can take both the low and high-resolution versions of the generated frame and force the super-resolution module to maintain consistency. Specifically, we take the ViT-based feature extractor networks of StyleGAN-t networks as is and run them two times, once on the low-resolution image and once on the high-resolution version of it. The feature extractors work at a constant resolution, we therefore bilinarly adjust the resolution as needed. Finally, the extracted features are concatenated along the channel axis and are fed to the trainable discriminator heads of StyleGAN-T. The generated video is passed through a video discriminator that has the same design as the video discriminator described in [4].
@@ 3.5 Implementation details
We utilize the most extensive version of the StyleGAN-T generator, which contains around one billion parameters, and apply its progressive training approach. Training starts with the generator producing $16\times 16$ resolution for both appearance and motion feature planes, and gradually, we increase this to $128\times 128$. The video output resolution remains unchanged during training, irrespective of the feature plane resolution. To match the increasing resolution of the feature planes, we progressively reduce the Gaussian blur applied to the frames of the dataset, starting from a sigma value of 10 and halving it each time the feature resolution is increased, following the sequence: $10,5,2.5,1.25,$ to 00. The transition to a higher resolution is triggered when the FID score ceases to improve for randomly sampled frames. The full training process takes about $5$ days on $8$ Nvidia A100 GPUs, with a consistent batch size of $128$ throughout the training sessions with Adam optimizer. We found that larger batch sizes have the effect of stabilizing training. For all evaluation purposes, we use the same learning rate for the generator($2.5\times 10^{-3}$) and discriminator($2\times 10^{-3}$) and Adam optimizer with $\beta_{1}=0$ and $\beta_{2}=0.99$. Following standard StyleGAN-T training, we use the R1 penalty for the discriminators. We use a loss weight ratio of $3:1$ between the video to image discriminator.
@@ 4 Experiments
We evaluate RAVEN on the video generation properties (Fig. 1), compare it to the state-of-the-art methods on synthetic and real datasets, and perform a detailed set of ablation studies.
@@ 4.1 Datasets
Our model was evaluated across various datasets featuring distinct motion types, specifically focusing on 1) Talking faces with all motions and a subset with 10 specific motions, 2) Fashion videos, and 3) UCF 101 dataset, all at a resolution of $256\times 256$. Similar to image-based GANs, we observed that video GANs also perform better with larger datasets. Due to the insufficiency of existing datasets in terms of size, we introduced the first two datasets.
Talking Faces:
Public datasets such as CelebV-HQ [46] and Face-Forensics [25] were inadequate in size and quality for our needs. We generated a synthetic dataset of approximately 400 thousand videos of talking faces using Thin-plate spline motion models [43] to animate FFHQ faces [18] with CelebV-HQ motions [46]. However, this process yielded many low-motion videos, prompting us to curate a subset of 10 high-motion videos from CelebV-HQ [46] to create a more dynamic Talking Faces dataset.
Fashion Videos:
To capture a variety of motions beyond facial movements, we created the Fashion Videos dataset. The original dataset from Dwnet [42] was too small, so we synthesized additional data by combining motions from pairs of videos. The Thin-plate spline [43] and First order motion models [26] proved inadequate for this task, leading us to adopt the Bidirectional Deformable Motion Modulation method [41]. We further refined the dataset by excluding videos with long dresses that caused inconsistencies, resulting in a total of 37,285 videos.
UCF101:
For the UCF101 dataset [29], we used the official data loader to extract 160-frame clips at 30 fps, with a 5-frame interval between clips. By merging training and test sets, we compiled a large dataset of 197,296 videos. We opted out of class conditioning, center-cropped all videos to the short side, and resized them to $256\times 256$ pixels using bilinear interpolation for training.
@@ 4.2 Comparisons
Following video generation research norms, we use Fréchet Video Distance (FVD) [31] and Fréchet Inception Distance (FID) [14] as our main evaluation metrics. Consistent with the approaches of Skorokhodov et al. [28], we utilize two variants of FVD—$\text{FVD}_{16}$ and $\text{FVD}_{128}$—which correspond to the analysis of video clips containing 16 and 128 frames, respectively. For FID assessment, we compare 50,000 randomly generated frames against 50,000 original frames sampled from the dataset. We adopt the FVD computation method from StyleGAN-V [28]. Notably, we opt for mp4 format to store generated and real videos to minimize memory usage. For instance, storing 2048 videos with 160 frames each consumes only 60 MB in mp4 format, in stark contrast to the 32 GB required for uncompressed png frames. This efficiency alleviates the memory constraints that previously limited StyleGAN-V [28] to using only 2048 samples for metric estimation. To maintain comparability with existing benchmarks, we continue to use 2048 samples for metric evaluation. Our method is benchmarked against the leading GAN-based approaches in unconditional video generation, namely StyleGAN-V [28] and MoCoGAN [30] with a StyleGAN2 backbone.
Figure 3: Qualitative results: Here we show the qualitative results generated by our method in comparison to MoCoGAN [30] and StyleGAN-V [28]. For visualization, we only show $6$ equally spaced frames from a video clip of length $5$ seconds.
Qualitative evaluation:
@@
In Fig. 3 RAVEN is qualitatively compared with StyleGAN-V [28] and MoCoGAN [30] by examining videos with a total duration of 5.3 seconds at 30 fps, selecting every 21st frame for display. The complete videos are provided in the supplementary materials in mp4 format. Our approach demonstrates its ability to capture the long-range spatial and temporal dependencies in the data, as shown by the consistency in the identity of subjects in talking face sequences and the coherent changes in appearance and limb movement in fashion videos (Fig. 1 and Fig. 3). Notably, in the talking faces dataset, StyleGAN-V shows a similar performance to our method, which could be due to two factors. Firstly, the model capacity required for representing talking faces, particularly with minimal motion, is well within StyleGAN2’s capabilities. Secondly, the dataset we synthesized from CelebV-HQ and FFHQ, using Thin-plate spline motion models, inherently features limited motion, particularly due to the prevalence of smiling expressions with partially open mouths in FFHQ, resulting in constrained jaw movement. These factors diminish the performance distinction between our method and StyleGAN-V. However, for the Talking Faces 10 motions dataset, which includes videos with more pronounced motions, our method shows a greater performance margin as indicated in the quantitative comparison table (Tab. 3). The superiority of our approach becomes more apparent in the fashion video dataset, which presents more complex structures such as articulated human figures, where other state-of-the-art methods struggle significantly.
@@
Talking Faces all/$10$ Motions Fashion Videos UCF 101 # T-FLOPs
@@
StyleGAN-V 11.8/18.6 135/135 112/138 19.66 167.3 167.9 1431 1773 10.125
MOCOGAN 58.89/– 209.4/– 348.2/– 202.9 686.6 1700 1630 2293 10.092
Ours-Stg2 323.3/– 4605/– 2781/– 298.3 967.0 2114 159.4 3011 6884 2.688
Ours-StgXL 26.49/– 116.4/– 203.0/– 26.28 72.46 219.2 81.6 1160 1982 7.686
Ours-StgT 8.9/4.7 97/65.6 110.3/62.1 10.8 67.5 111.1 79.7 1136 1933 4.183
Table 3: We compare the performance of our method against state-of-the-art diffusion-based and GAN-based methods under three different datasets and two different metrics, FID and FVD. We use $2048$ video samples, $16$ and $128$ consecutive frames to estimate FVD. To estimate FID, we use real random frames from all the videos available and 50K generated random video frames.
Our evaluation, detailed in the Tab. 3, indicates that our video generation method outperforms current leading GAN-based alternatives. This success is attributed to our method’s ability to train across the entire video volume, unlike other state-of-the-art (SOTA) approaches. Our model also demonstrates a significant reduction in computational effort to generate a single video sample of 160 frames, requiring less than half the FLOPs compared to other SOTA models. However, our use of a per-frame super-resolution network, while parallelizable and rapid (running at 54 FPS on a Quadro RTX 6000 GPU), does increase the total computational expense, accounting for about 50% of our compute budget.
@@
The FID and FVD metrics on the UCF101 dataset [29] are higher for all models (Tab. 3), likely due to the dataset’s intricate motion patterns and the broad spectrum of action categories it contains. Moreover, the limited sample size used for evaluation ( 2K videos) may not suffice for an accurate quality assessment, especially highlighted by the $\text{FVD}_{128}$ metric.
@@
Our findings show that while our model achieves competitive results at a resolution of $256\times 256$ with a relatively low FID, the task of generating high-resolution videos remains a challenge. Some models exhibit strengths at lower resolutions but experience performance decline at higher resolutions. This highlights the dynamic nature of video generation research, where each approach has its advantages and limitations, and where there is a continuous search for methods that balance performance across various resolutions and complexities.
@@ 4.3 Ablation Studies
We also identify that the dynamic nature of video content, which frequently involves the translation of objects or entire scenes, can lead to redundant feature replication that could otherwise be utilized for enhancing representational capacity. To address this, we introduce a novel concept of feature flow within our model, designed to optimize the utilization of representational capacity by minimizing redundancy. The effectiveness of this design is evident in the improved performance metrics ( Tabs. 2 and 1).
In addition, we employ the most expansive version of StyleGAN currently available, which we posit as the most suitable choice for our framework. This assertion is substantiated by a comparative analysis against a lesser model employing StyleGAN2’s default settings. The comparative results, reflected in the rows labeled Ours-Stg2, Ours-StgXL, and Ours-Full in our quantitative comparison table, clearly demonstrate the advantage of utilizing the larger StyleGAN model in our approach. In Ours-Stg2, Ours-StgXL, and Ours-Full we have used StyleGAN-2, StyleGAN-XL, and StyleGAN-t as the backbone of our generator respectively. The number of trainable parameters increases significantly from StyleGAN-2 to StyleGAN-XL and from StyleGAN-XL to StyleGAN-t. This selection is not arbitrary but is justified by the superior performance and representational capabilities it provides for our video generation framework.

### v2

@@ 3.3 CNN backbone of Tri-plane and generation
The tri-plane’s composition of three 2D feature arrays allows for the use of a conventional 2D CNN generator. We selected the StyleGAN-T architecture for its exceptional capacity for feature representation. Adapting StyleGAN-T’s output, our model generates a $64\times 64\times 68$ feature grid, partitioned into a 32-channel global feature grid ($F_{G}$) and three $12$-channel motion feature grids. From these, we derive local and global flow fields and a mask, which are utilized to warp and blend the appearance features, forming an appearance feature volume. To farther enhance training efficiency, we sample a frame at a random time $t$ at a $64\times 64$ resolution and generate the entire video at $32\times 32\times 32$. We achieve the final $256\times 256$ resolution through image-domain super-resolution, applied only to the selected frame to conserve computational resources. This process mirrors the approach taken in eg3D, utilizing two StyleGAN2 synthesis blocks to upscale the image by a factor of four.
@@ 3.4 Double Discrimination
Similar to a standard GAN setting, we use a discriminator to criticize the generated data. We deploy two different discriminators, one for the frame of time step $t$, and the other for the generated video. We extend the built-in StyleGAN-T discriminator to accommodate $6$ channels instead of $3$. Such that it can take both the low and high-resolution versions of the generated frame and force the super-resolution module to maintain consistency. Specifically, we take the ViT-based feature extractor networks of StyleGAN-t networks as is and run them two times, once on the low-resolution image and once on the high-resolution version of it. The feature extractors work at a constant resolution, we therefore adjust the resolution as needed with bilinear interpolation. Finally, the extracted features are concatenated along the channel axis and are fed to the trainable discriminator heads of StyleGAN-T. The generated video is passed through a video discriminator that has the same design as the video discriminator described in [6].
@@ 3.5 Implementation details
We utilize the most extensive version of the StyleGAN-T generator, which contains around one billion parameters, and apply its progressive training approach. Training starts with the generator producing $16\times 16$ resolution for both appearance and motion feature planes, and gradually, we increase this to $128\times 128$. The video output resolution remains unchanged during training, irrespective of the feature plane resolution. To match the increasing resolution of the feature planes, we progressively reduce the Gaussian blur applied to the frames of the dataset, starting from a sigma value of 10 and halving it each time the feature resolution is increased, following the sequence: $10,$ $5,$ $2.5,$ $1.25,$ 00. The transition to a higher resolution is triggered when the FID score ceases to improve for randomly sampled frames. The full training process takes about $5$ days on $8$ Nvidia A100 GPUs, with a consistent batch size of $128$ throughout the training sessions with Adam Optimizer. We found that larger batch sizes have the effect of stabilizing training. For all evaluation purposes, we use the same learning rate for the generator($2.5\times 10^{-3}$) and the discriminator($2\times 10^{-3}$) and Adam optimizer with $\beta_{1}=0$ and $\beta_{2}=0.99$. Following standard StyleGAN-T training, we use the R1 penalty for the discriminators. We use a loss weight ratio of $3:1$ between the video to image discriminator.
@@ 4 Experiments
We evaluate RAVEN on the video generation properties (Fig. 1), compare it to the state-of-the-art methods on synthetic and real datasets, and perform a detailed set of ablation studies. Here we primarily focus on comparing our method to other GAN-based methods, keeping in mind apples-to-apples comparison. We however acknowledge the importance of application-focused comparisons, i.e. comparison against all works that generate videos, such as [23, 22, 53]etc. We separate this in Sec. 4.4.
@@ 4.1 Datasets
Our model was evaluated across various datasets featuring distinct motion types, specifically focusing on 1) Talking faces with all motions and a subset with 10 specific motions, 2) Fashion videos, and 3) WebVid10M-Flowers dataset, all at a resolution of $256\times 256$. Similar to image-based GANs, we observed that video GANs perform better with larger datasets. Due to the insufficiency of existing datasets in terms of size, we introduced the first two datasets. Additionally, we collected videos of flowers moving with the wind to form a subset of the WebVid10M [2] dataset to evaluate our model on an existing real-world dataset. According to the download page, the original WebVid10M dataset is no longer available.
Talking Faces: Public datasets such as CelebV-HQ [55] and Face-Forensics [32] are inadequate in size and quality for our needs. We generated a synthetic dataset of approximately 400 thousand videos of talking faces using Thin-plate spline motion models [52] to animate FFHQ faces [24] with CelebV-HQ motions [55]. However, this process yielded many low-motion videos, prompting us to curate a subset of 10 high-motion videos from CelebV-HQ [55] to create a more dynamic Talking Faces dataset.
Fashion Videos: To capture a variety of motions beyond facial movements, we created the Fashion Videos dataset. The original dataset from Dwnet [51] is too small, so we synthesize additional data by combining motions from pairs of videos. The Thin-plate spline [52] and First order motion models [35] proved inadequate for this task, leading us to adopt the Bidirectional Deformable Motion Modulation method [50]. We further refined the dataset by excluding videos with long dresses that caused inconsistencies, resulting in a total of 37,285 videos.
WebVid10M-Flowers: To form the WebVid10M-Flowers dataset we collect a subset of videos from WebVid10M [2] dataset that contain flowers moving in the wind. We perform this using similarity matching of the captions of the videos. We use the sentence transformer [31] to perform sentence-level embedding and retrieval. The query that we used is “flowers moving in the wind”.
@@ 4.2 Comparisons
Following video generation research norms, we use Fréchet Video Distance (FVD) [39] and Fréchet Inception Distance (FID) [19] as our main evaluation metrics. Consistent with the approaches of Skorokhodov et al. [37], we utilize two variants of FVD, namely, $\text{FVD}_{16}$ and $\text{FVD}_{128}$—which correspond to the analysis of video clips containing 16 and 128 frames, respectively. For FID assessment, we compare 50,000 randomly generated frames against 50,000 original frames sampled from the dataset. We adopt the FVD computation method from StyleGAN-V [37]. Notably, we opt for mp4 format to store generated and real videos to minimize memory usage. For instance, storing 2048 videos with 160 frames each consumes only 60 MB in mp4 format, in stark contrast to the 32 GB required for uncompressed png frames. This efficiency alleviates the memory constraints that previously limited StyleGAN-V [37] to using only 2048 samples for metric estimation. To maintain comparability with existing benchmarks, we continue to use 2048 samples for metric evaluation. Our method is benchmarked against the leading GAN-based approaches in unconditional video generation, namely StyleGAN-V [37] and MoCoGAN [38] with a StyleGAN2 backbone. We also compare with the state-of-the-art open source diffusion-based video generation method ‘stable video diffusion (SVD)’ [4]. We note that strictly speaking it is not a fair comparison as the model is image conditioned i.e., it takes as input the first frame. Furthermore, to compare under the same dataset we finetune a pre-trained SVD model on our dataset. Therefore it has an overwhelming data advantage over any of the GAN-based models. We still include these results in our quantitative evaluation section.
@@
Qualitative evaluation: In Fig. 3(a) and 3(b) RAVEN is qualitatively compared with StyleGAN-V [37] and MoCoGAN [38] with enhanced backbones as well as stable video diffusion (SVD) [4]. Here we examine videos with a total duration of $5.3$ seconds at $30$ fps by selecting every $21$st frame for display. However, since SVD only generates 25 frames, we simply display 6 equally spaced frames. The complete videos are provided in the supplementary materials in mp4 format. In  (Fig. 1) by overlapping different color channels, we highlight the movement of the objects generated by our method. Moreover, our approach demonstrates its ability to capture the long-range spatial and temporal dependencies in the data, as shown by the consistency in the identity of subjects in talking face sequences and the coherent changes in appearance and limb movement in fashion videos. Notably, in the talking-faces dataset, StyleGAN-V shows a similar performance to our method, which could be due to two factors. Firstly, the model capacity required for representing talking faces, particularly with minimal motion, is well within StyleGAN2’s capabilities. Secondly, the dataset we synthesized from CelebV-HQ and FFHQ, using Thin-plate spline motion models, inherently features limited motion, particularly due to the prevalence of smiling expressions with partially open mouths in FFHQ, resulting in constrained jaw movement. These factors diminish the performance distinction between our method and StyleGAN-V. However, for the Talking Faces 10 motions dataset, which includes videos with more pronounced motions, our method shows a greater performance margin as indicated in the quantitative comparison table (Tab. 3). The superiority of our approach becomes more apparent in the fashion video dataset, which presents more complex structures such as articulated human figures, where other state-of-the-art methods struggle significantly.
Figure 3: Qualitative results: Here in Fig. 3(a) and Fig. 3(b) we show the qualitative results generated by our method in comparison to MoCoGAN [38], StyleGAN-V [37] and stable video diffusion (SVD) [4]. For visualization, we only show $6$ equally spaced frames from a video clip of length $5$ seconds for all models except for SVD. Since SVD only generates 25 frames, we simply visualize 6 equally spaced frames (i.e. we skip 4 frames in between every shown frame for SVD). We provide several generated videos from each model in the supplementary material.
@@
Talking Faces all/$10$ Motions Fashion Videos WebVid10M-Flowers # T-FLOPs
@@
StyleGAN-V 11.8/18.6 135/135 112/138 19.66 167.3 167.9 32.30 142.2 10.125
StyleGNA-V-T 166.5/– 257.6 2851 302.4 1259 1532 303.7 744.6 43.039
MOCOGAN 58.89/– 209.4/– 348.2/– 202.9 686.6 1700 38.64 145.9 10.092
Ours-Stg2 323.3/– 4605/– 2781/– 298.3 967.0 2114 FT FT FT 2.688
Ours-StgXL 26.49/– 116.4/– 203.0/– 26.28 72.46 219.2 FT FT FT 7.686
Ours-StgT 8.9/4.7 97/65.6 110.3/62.1 10.8 67.5 111.1 8.71 119 190 4.183
Table 3: We compare the performance of our method against state-of-the-art diffusion-based and GAN-based methods under three different datasets and two different metrics, FID and FVD. We use $2048$ video samples, $16$, and $128$ consecutive frames to estimate FVD. To estimate FID, we use real random frames from all the videos available and 50K generated random video frames. FT: fails to train.
Our evaluation, detailed in the Tab. 3, indicates that our video generation method outperforms current leading GAN-based alternatives [^1] We forego comparison with MOCOGAN-HD with its original backbone as its performance is comprehensively beaten by StyleGAN-V, and we outperform StyleGNA-V on all datasets under all metrics. We instead compare to the enhanced version of MOCOGAN proposed in StyleGAN-V.. This success is attributed to our method’s ability to train across the entire video volume, unlike other state-of-the-art (SOTA) approaches. We investigate if StyleGAN-V’s performance can be improved by simply changing the generative backbone to a stronger model as compared to StyleGAN2, namely StyleGAN-T. We call this method StyleGAN-V-T. However, as noted in [33], we also observe that simply changing the generative backbone is not enough and in fact it degrades its performance. Our model also demonstrates a significant reduction in computational effort to generate a single video sample of $160$ frames, requiring less than half the FLOPs compared to other SOTA models. However, our use of a per-frame super-resolution network, while parallelizable and rapid (running at 54 FPS on a Quadro RTX 6000 GPU), it does increase the total computational expense, accounting for about 50% of our compute budget at test time.
Furtheremore, we compare under the same metric the performance of SVD [4] against our method. We fine-tune a pre-trained SVD model on our datasets for this comparison. However, as mentioned earlier there remains a lot of inherent nonuniformity in training procedures simply because the two methods are widely different in their scope. First, SVD generates only 25 frames as compared to 160 frames that our method generates. This immediately poses the challenge of framerate selection for the training dataset. We subsample our training dataset by 5 fold to preserve a similar amount of motion complexity for SVD. Second, SVD uses the first frame as a condition, therefore FID computation on the first frame becomes noninformative. We therefore do not report this metric. Third, the FVD2048_128f metric requires 128 frames in every sample. This is unavailable for SVD, we therefore report only FVD2048_16f. The performance of SVD under the FVD2048_16f metric are $87.7$, $260.9$, and $196.6$ for the three of our datasets talking-faces, fashion videos, WebVid10M-Flowers respectively. Please refer to the Supmat for qualitative results. As can be seen, RAVEN is competitive to SVD despite SVD having been pretrained on data that is many-fold larger than ours.
@@
Our findings show that while our model achieves competitive results at a resolution of $256\times 256$ with a relatively good FID score, the task of generating high-resolution videos remains a challenge. Some models exhibit strengths at lower resolutions but experience performance decline at higher resolutions. This highlights the dynamic nature of video generation research, where each approach has its advantages and limitations, and where there is a continuous search for methods that balance performance across various resolutions and complexities.
@@ 4.3 Ablation Studies
We also identify that the dynamic nature of video content, which frequently involves the translation of objects or of entire scenes, can lead to redundant feature replication that could otherwise be utilized for enhancing representational capacity. To address this, we introduce the novel concept of feature flow within our model, designed to optimize the utilization of representational capacity by minimizing redundancy. The effectiveness of this design is evident in the improved performance metrics (Tabs. 2 and 1).
In addition, we employ the most expressive version of StyleGAN currently available, which we posit as the most suitable choice for our framework. This assertion is substantiated by a comparative analysis against a lesser model employing StyleGAN2’s default settings. The comparative results, reflected in the rows labeled Ours-Stg2, Ours-StgXL, and Ours-Full in Tab. 1, clearly demonstrate the advantage of utilizing the larger StyleGAN model in our approach. In Ours-Stg2, Ours-StgXL, and Ours-Full we have used StyleGAN-2, StyleGAN-XL, and StyleGAN-t as the backbone of our generator respectively.
4.4 Comparisons to Non-GAN-based approaches
We summarize the performance of recent video generative models from diverse backgrounds and design choices in Sec. 4.2. We report FVDs as are found in the published papers of the indicated methods. There are variations in the training and evaluation protocol, such as crop size, normalization scheme, compression used to store original data, feature extraction network used etc. as also noted in [37]. This makes drawing a quick conclusion a hard task as the precise FVD number is sensitive to the factors mentioned above. We however do think that one obtains useful insight from this comparison nonetheless. Here we would like to highlight, as an example, the comparison against Vidoe-LDM [5]. It achieves an FVD16 of $550.6$ in UCF 101 dataset at a resolution $128\times 256$ which is 2×2$\times$2 × lower than us. Despite being a diffusion-based model as opposed to our GAN based model, it uses a similar number of model parameters. However, it requires several stages of denoising and therefore significantly more floating point operations.